OpenStudy (anonymous):

If mean of number is 4.5... what is the sum of numbers...?

OpenStudy (anonymous):

The mean of a set of numbers is equal to: \[\frac{Sum.of.numbers}{\#.of.numbers}\]

OpenStudy (anonymous):

The sum will be equal to the number of values times the mean.

OpenStudy (johnweldon1993):

Right..so if it tells you how many numbers were added to make said "mean" ...you need only multiply the mean times the amount of numbers

OpenStudy (anonymous):

please solve here

OpenStudy (anonymous):

@ johnweldon.... i did this simple step, but i think its wrong.... 4.5*1=4.5 sum of numbers...

OpenStudy (anonymous):

How many numbers, or sets of data is being used to calculate the mean. You can't find the sum without that.

OpenStudy (anonymous):

@ heril... yeah dear... in question i have only the mean of the number.... and need the sum... please help...

OpenStudy (anonymous):

Without the information of the size of the set, survey, group, whatever you are trying to find the mean of, you cannot find the sum of the set from the mean. Is there any additional information in the instructions?

OpenStudy (johnweldon1993):

Right @engrbill can you type the whole question please? Or is that everything?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that is the full question...

OpenStudy (johnweldon1993):

If that is the full question...we can have multiple answers... \[\large \frac{1 + 7 + 2 + 8}{4} = 4.5\] That sum would be 18... But we could also have \[\large \frac{1 + 2 + 3 + 5 + 7 + 9}{6} = 4.5\] and that sum would be 27 So it is impossible to say...we do not know how many numbers you are given to calculate that mean...and therefor we will have an endless amount of answers.

OpenStudy (anonymous):

ok, bundle of thanks